OpenStudy (anonymous):

For the simple harmonic motion equation d=8sin(4pi t) , what is the period?

OpenStudy (blockcolder):

For \(y=a \sin(bx)\), the period is \(\Large \frac{2\pi}{b}\).

OpenStudy (blockcolder):

In this case, \(b=4\pi\) so the period is \(\Large\frac{2\pi}{4\pi}=\frac{1}{2}\) units.

OpenStudy (anonymous):

ty

OpenStudy (blockcolder):

No prob.